OpenStudy (anonymous):

Solve the system by either substituition or elimination: 2x+5y=-22 5x-7y=23

OpenStudy (anonymous):

using elimination method it will be[5(2x+5y=-22)]which is 10x+25y=110 as the first equation.the second equation will be[2(5x-7y=23)]. 10x-14y=46. subtracting the first equation from the second one it will be 39y=64.divide throughout by 39 to find value of y. y=64/39.find value of x by substituting the value of y in either second or first equation. 10x+25(64/39)=110. x=[110-25(64/39)]/10.as simple as that.